OpenStudy (anonymous):

y= 4.75 sin (3x-1) -5 max= min= amp= p= m= sp=

OpenStudy (anonymous):

max value sine =1 so max y = 4.75 - 5 = -0.25 min = -4.75-5 = -9.75

OpenStudy (anonymous):

and what about the amp, p, m, and sp or start point

OpenStudy (anonymous):

???